# CHIRLA: Comprehensive High-resolution Identification and Re-identification for Large-scale Analysis

### v1

@@ Abstract
Person re-identification (Re-ID) is a key challenge in computer vision, requiring the matching of individuals across different cameras, locations, and time periods. While most research focuses on short-term scenarios with minimal appearance changes, real-world applications demand robust Re-ID systems capable of handling long-term scenarios, where persons’ appearances can change significantly due to variations in clothing and physical characteristics. In this paper, we present CHIRLA, Comprehensive High-resolution Identification and Re-identification for Large-scale Analysis, a novel dataset specifically designed for long-term person Re-ID. CHIRLA consists of recordings from strategically placed cameras over a seven-month period, capturing significant variations in both temporal and appearance attributes, including controlled changes in participants’ clothing and physical features. The dataset includes 22 individuals, four connected indoor environments, and seven cameras. We collected more than five hours of video that we semi-automatically labeled to generate around one million bounding boxes with identity annotations. By introducing this comprehensive benchmark, we aim to facilitate the development and evaluation of Re-ID algorithms that can reliably perform in challenging, long-term real-world scenarios.
@@
Person re-identification (Re-ID) is a challenging problem in computer vision that involves matching individuals across different camera views, locations and times. This task becomes even more complex when considering long-term scenarios, where persons may change their appearance due to variations in clothing or alterations in physical characteristics.
Early Re-ID research relied on datasets that primarily focused on short-term Re-ID, involving single images or small image sequences. Examples include ETH [1], GRID [2], VIPeR [3], and CAVIAR [4]. These datasets typically feature low-resolution images and limited temporal information. In later years, with the advent of deep learning, larger datasets began to appear. These datasets expanded the number of identities (IDs), provided more views of the same person from different cameras and offered longer sequences. This is the case of datasets such as CUHK03 [5], RAiD [6] and Market1501 [7].
The most recent datasets directly provide the complete videos of the sequences, with the necessary bounding boxes and ID annotations, so that they can be used to evaluate the actual performance of video re-identification systems, and can take advantage of complementary tracking techniques. In this category we can find the datasets MARS [8] and PRW [9].
While many of these existing datasets focus on short-term Re-ID, there is a growing need for datasets that capture the variability of individuals’ appearances over extended periods. Such datasets are crucial for developing more robust Re-ID algorithms capable of handling real-world challenges, particularly in environments where individuals frequently change their clothing or physical characteristics. Existing datasets often lack the variability in clothing and appearance changes, making them less suitable for training and evaluating models designed for long-term Re-ID scenarios.
Table 1 summarizes the characteristics of the datasets mentioned above. In this table, time coherence refers to whether the cameras in the analyzed dataset recorded simultaneously and were synchronized across views, enabling the identification of the same person across different cameras or during transitions between camera viewpoints in the environment setup. This feature also includes the requirement of providing the corresponding videos from that setup, allowing their analysis for person Re-ID tasks. The listed resolution corresponds to the maximum resolution available in each dataset. In cases where cameras or sequences were recorded at different resolutions, only the highest resolution is reported. Some datasets provide cropped ID bounding boxes directly, while others offer full-frame images. For the CUHK03 dataset [5], the resolution refers to bounding box crops, which are either manually created or automatically detected, with resolutions varying around 160×60 pixels.
@@
/Videos (s)
# IDs
Total
Frames
@@
Cameras
Time
Coherence
IDs
Coherence
ETH [1] 4 38 146 2,293 640×480 1 x x
GRID [2] - - 250 1,275 320×230 8 x
VIPeR [3] - - 632 - 128×48 2 x
CAVIAR [4] - - 72 1,220 72×144 2 x
CUHK03 [5] - - 1,360 13,164 Variable* 6 x
RAiD [6] - - 43 6,920 128×64 4 x
Market1501 [7] - - 1,501 32,668 1080×1920 6 x
MARS [8] $\sim$20,000 2 1,261 1,191,003 1080×1920 6 x
PRW [9] - - 932 11,816 1080×1920 6 x
CHIRLA 70 284 22 596,345 1080×720 7
Table 1: Summary of the reviewed datasets and their features. Time coherence indicates whether cameras recorded simultaneously and were synchronized, enabling person re-ID analysis across viewpoints. The listed resolution corresponds to the maximum resolution available; for datasets with varying resolutions, only the highest is reported.
Although many of these datasets include a large number of IDs, they are often constrained by the short durations of their recordings, limiting their utility for long-term Re-ID evaluations. In most cases, once an individual appears in one video or sequence, they are not annotated or re-identified in subsequent sequences, restricting the evaluation of Re-ID algorithms designed for long-term scenarios.
This paper introduces a novel dataset specifically designed for long-term person (Re-ID) in an indoor environment. The dataset captures a multi-room office space featuring interconnected indoor areas where movement follows realistic patterns, including hallways, shared workspaces, and entry points. Recordings were obtained from multiple cameras strategically placed around the office, covering various angles and perspectives. Data collection was conducted over a seven-month period to capture the natural temporal variations in participants’ appearances. To further enhance the complexity of the dataset, participants were instructed to change their clothing between recording sessions, introducing significant variability in appearance that challenges existing Re-ID algorithms. The primary objective of this dataset is to provide a comprehensive benchmark for evaluating Re-ID models in real-world, long-term scenarios where individuals’ clothing and physical appearance may vary significantly over time. By combining diverse recordings with controlled variations in appearance, this dataset offers a valuable resource for advancing research in person Re-ID, particularly for applications requiring high reliability over extended periods.
@@ Methods
This study was conducted under the supervision of the Ethical Committee of the University of Alicante (Approval No. UA-2022-11-12). All subjects were informed about the study, including how their data would be handled, and provided their consent to participate and share the collected data. Specifically, participants granted permission for their inclusion in dataset video recordings, which capture physical data (face and body features in video recordings), and consented to making these recordings publicly available. The dataset has been made publicly available under the CC-BY license for academic use only and is not permitted for commercial purposes.
@@
To streamline and expedite the manual verification and correction of bounding boxes and ID annotations, we developed a graphical user interface (GUI). This tool significantly facilitates the labeling process (Figure 3). The GUI is publicly available at [13].
@@
$FN$ ={}={}= total number of false negatives over all frames.
$FP$ ={}={}= total number of false positives over all frames.
$IDSW$ ={}={}= total number of ID switches over all frames.
$GT$ ={}={}= total number of ground truth detections over all frames.
@@
$Acc_{k}=\left\{\begin{array}[]{ll}1,&\text{if top-$k$ ranked gallery samples contain the query ID}\\ \\ 0,&\text{otherwise}\end{array}\right.$ (6)
@@
By default, precision takes all the retrieved correspondences into account, but it can also be evaluated at a given number of retrieved images, commonly known as cut-off rank, where the model is only assessed by considering its top-most queries. The measure is called precision at $k$ or P@$k$. Using the P@$k$ we can calculate the average precision (AP@$n$) for a number $n$ of matches, as shown in Equation 8. The final AP is the value of AP@$n$ when all the ground truth positives have been obtained.
@@
$GTP$ ={}={}= total number of ground truth positives (correct samples of the query class),
n ={}={}= total number of candidate matches ($n$ elements most similar to the query),
$rel@k$ ={}={}= relevance function that equals 1 if the ID at rank $k$ is correct.
@@
$N$ ={}={}= total number of queries,
$i$ ={}={}= index for the i-th query,
@@
The dataset is composed of a collection of videos, as provided by the set of cameras in the environment, and the corresponding Re-ID labels. The labels comprise the bounding box of the individuals present in each frame of the video and the corresponding ID. In the following Subsections, we will delve in the components of the dataset and the organization details.
@@
The dataset consists of 10 sequences, each containing seven videos recorded by the seven cameras deployed in the environment. The videos were collected over a period of seven months, allowing the dataset to capture variations in both the environment and the appearance of individuals. The average duration of a single video is 284 seconds, with a range from 113 to 567 seconds. In total, the dataset comprises more than 5.52 hours of video and 596,345 frames, with a total of 963,554 annotated bounding boxes. The details of each sequence are presented in Table 2. The videos are stored in AVI format, encoded using DivX MPEG-4 at 30 fps with a bitrate of 4048 Kbps.
@@
In addition to the video data, we provide Re-ID labels. As mentioned earlier, each video is accompanied by a label file that contains, for each frame, the bounding boxes of all detected individuals along with their corresponding IDs. The labels are stored in JSON format, making them easy to parse. Each JSON file consists of a dictionary where the keys represent the frame numbers of the video, and the corresponding values are lists. Each element in these lists is a dictionary containing two keys: "id", which stores the unique ID assigned to each individual, and "BboxP", which contains a list of the bounding box coordinates $[X_{1},Y_{1},X_{2},Y_{2}]$ that define the region occupied by the person in that frame. Frame and ID values are 1-based, whereas bounding box coordinates are 0-based. Specifically, the top-left corner of the bounding box corresponds to $(0,0)$. Figure 4 presents examples of labeled individuals in randomly selected frames from the dataset.
@@
The dataset is organized into two primary directories: annotations and videos. Each directory contains 10 subdirectories corresponding to the individual sequences listed in Table 2. These sequences are labeled as seq_XXX, where XXX represents a three-digit, zero-padded sequence number (e.g., seq_026). Within each sequence subdirectory, the data is further divided by camera views, with seven files provided per sequence, one for each camera view. Each file is named based on the camera number, as well as the date and time when the videos were recorded.
@@
These tests assess the ability of a system to detect a person in the scene, create an ID for them, and track them on a single camera until the person disappears from the scene for a few seconds. For this purpose, the video sequences are divided into a training set and a test set. The training set is used to optimize certain parameters of the tracking algorithm that make it applicable to this setting, and adapt them to the resolution and the distance at which persons are seen in the scene. On the other hand, the test set will be used to evaluate the performance of the tracking systems, making use of the MOTA, IDF1 and their auxiliary metrics, as well as the generalization capacity of these systems after having optimized their parameters in videos from training sequences only.
With the proposed training and test sets, we check the capabilities of the tracking system in the following scenarios:
@@
For this purpose, the video sequences are divided into training and test sets. The training sets are used to optimize the parameters of the Re-ID algorithm, adapting it to the visual characteristics of the recorded persons, so that it learns to recognize their ID, especially focusing on the characteristics that allow them to be identified in the long term.
On the other hand, the test sets are designed to evaluate the performance of the system in different Re-ID scenarios, using metrics such as mAP, CMC and other auxiliary metrics that allow measuring its generalization capacity after the optimization of parameters in sequences other than training ones.
With the proposed training and test sets, we check the capabilities of the Re-ID system in the following scenarios:

### v2

@@ Abstract
Person re-identification (Re-ID) is a key challenge in computer vision, requiring the matching of individuals across cameras, locations, and time. While most research focuses on short-term scenarios with minimal appearance changes, real-world applications demand robust systems that handle long-term variations caused by clothing and physical changes. We present CHIRLA, Comprehensive High-resolution Identification and Re-identification for Large-scale Analysis, a novel dataset designed for video-based long-term person Re-ID. CHIRLA was recorded over seven months in four connected indoor environments using seven strategically placed cameras, capturing realistic movements with substantial clothing and appearance variability. The dataset includes 22 individuals, more than five hours of video, and about 1M bounding boxes with identity annotations obtained through semi-automatic labeling. We also define benchmark protocols for person tracking and Re-ID, covering diverse and challenging scenarios such as occlusion, reappearance, and multi-camera conditions. By introducing this comprehensive benchmark, we aim to facilitate the development and evaluation of Re-ID algorithms that can reliably perform in challenging, long-term real-world scenarios. The benchmark code is publicly available at: https://github.com/bdager/CHIRLA.
@@
Person re-identification (Re-ID) is a challenging problem in computer vision that involves matching individuals across different camera views, locations and times. This task becomes even more complex when considering long-term scenarios, where individuals may change their appearance due to variations in clothing or alterations in physical characteristics.
Early Re-ID benchmarks such as ETH [1], GRID [2], VIPeR [3], and CAVIAR [4] focused on short-term conditions, typically with low-resolution images and limited temporal information. With the advent of deep learning, larger datasets such as CUHK03 [5], RAiD [6], Market-1501 [7], PRW [8], DukeMTMC [9] and MSMT17 [10] expanded scale and camera coverage, offering more identities (IDs) and additional views of the same person across cameras. However, these datasets remain fundamentally short-term: once an individual appears in one sequence, they are rarely annotated across different sessions, which limits their utility for long-term evaluation.
Video-based datasets such as MARS [11] advanced the field by providing video sequences with bounding box annotations, enabling temporal modeling and integration with tracking methods. Nevertheless, it remains short-term, restricted to recordings of only a few seconds per ID, and was not designed to capture deliberate appearance variation such as clothing changes.
Clothing-change benchmarks were later introduced to capture long-term variability. Motion-ReID [12] emerged as a video-based person Re-ID dataset, but it is constrained to very limited appearance changes, providing only two different outfits per ID recorded from two indoor disjoint static surveillance cameras.
Other datasets attempted to address clothing-change variability, but they are mainly image-based. Celeb-reID [13] consists of images collected from Internet street snapshots of celebrities. It is valuable for studying clothing variation and provides diverse appearance changes. However, the images are sourced from photographers’ snapshots, which do not replicate the camera angles, viewpoints, or environmental conditions typical of surveillance systems. PRCC [14] contains individuals captured by three cameras under two outfits per individual. Its paired design makes it suitable for controlled clothing-change evaluation, but it is limited in the extent of appearance variation, constrained to a small number of viewpoints, and remains image-based. LTCC [15] expands the setting by including more cameras and recording across multiple days with clothing labels, thereby increasing viewpoint coverage. Yet, like Celeb-reID and PRCC, LTCC consists of independent images without temporal continuity.
DeepChange [16] represents the largest dataset targeting long-term Re-ID under clothing change to date, collected over 12 months from 17 outdoor surveillance cameras. It provides valuable large-scale coverage of appearance variation. However, while the data originates from real-world video recordings, the public release consists only of low-resolution, fixed-size cropped bounding-box images rather than continuous tracklets or videos. This format, while realistic for outdoor surveillance, restricts evaluation primarily to body-based features and makes it difficult to assess Re-ID methods that also incorporate facial cues, which are particularly relevant in indoor scenarios. It also limits the dataset’s direct use of methods that can combine and take advantage of tracking or occlusion-aware benchmarks.
Table 1 summarizes the characteristics of the datasets discussed above. Overall, to the best of our knowledge, there is no existing person Re-ID dataset that is both long-term and video-based, offering temporal continuity, clothing variation, and image quality suitable for indoor surveillance analysis.
@@
/Video (s)
# IDs # Bboxes
Avg.
#Bboxes/ID
@@
Cam.
$C$ TC
ETH [1] 4 38 146 8580 59 640×480 1
GRID [2] - - 250 1,275 5 320×230 8
VIPeR [3] - - 632 1,264 2 128×48 2
CAVIAR [4] - - 72 1,220 17 72×144 2
CUHK03 [5] - - 1,360 13,164 10 Variable 6
RAiD [6] - - 43 6,920 161 128×64 4
Market1501 [7] - - 1,501 32,668 22 1080×1920 6
PRW [8] - - 932 34,304 37 1080×1920 6
DukeMTMC [9] - - 1,812 36,411 20 1920×1080 8
MSMT17 [10] - - 4,101 126,441 31 Variable 15
MARS [11] 20,715 2 1,261 1,067,516 847 1080×1920 6
Motion-ReID [12] 240 3.3 30 24,480 816 Variable 2
Celeb-reID [13] - - 590 10,842 18 128×64 1
PRCC [14] - - 221 33,698 152 Variable 3
LTCC [15] - - 152 17,138 113 Variable 12
DeepChange [16] - - 1,121 178,407 159 1920×1080 17
CHIRLA 70 284 22 963,554 43,798 1080×720 7
Table 1: Summary of the reviewed datasets and their main characteristics. $C$: clothing-change; TC: temporal coherence, indicating that cameras recorded simultaneously and were synchronized across views; and the recorded videos are provided to allow analysis of cross-camera person Re-ID. Image resolution corresponds to the maximum recording resolution available in each dataset; if multiple camera resolutions are used, only the highest is reported. For datasets that provide only cropped bounding boxes, the bounding-box resolution is reported. When marked as Variable, crops are either automatically detected by a person detector or annotated without fixed size constraints.
To address these gaps, we introduce CHIRLA, a novel dataset for long-term person Re-ID in indoor surveillance environments. CHIRLA was recorded over seven months (May–December) in a multi-room office space, capturing realistic movement patterns across hallways, shared workspaces, and entry points. Recordings were obtained from multiple cameras strategically placed around the office, covering various angles and perspectives, often ensuring visible faces in addition to body appearance. With this setup, CHIRLA includes variable resolutions depending on camera placement: lower-resolution crops in distant hallway views and higher-resolution crops in closer office settings. To further enhance the complexity of the dataset, participants were instructed to change their clothing between sessions, introducing significant appearance variability over time.
CHIRLA dataset introduces several unique features: (1) More bounding boxes ($\sim 964K$) and more samples per ID ($\sim 44K$) than previous clothing-change datasets, enabling deeper temporal analysis of the same individuals. (2) Continuous video sequences with detailed annotations, explicitly capturing occlusions (partial and complete), reappearances, and multi-camera transitions. (3) Fine-grained benchmark protocols for both Re-ID and tracking under occlusion, reappearance, and viewpoint changes. (4) Open-set evaluation with distractors: known IDs appear in both training and testing, while unknown IDs (distractors) appear only in testing, forcing models to set realistic decision thresholds.
In summary, CHIRLA complements existing datasets by providing temporally grounded and scenario-specific evaluations, while also enabling face-aware analysis thanks to the higher resolution of its person bounding boxes. Although it contains fewer IDs than large-scale outdoor benchmarks, CHIRLA offers richer per-ID data, an indoor surveillance setup, and realistic protocols for long-term Re-ID and tracking applications. CHIRLA is not designed as a large-scale training dataset, but rather as a complementary benchmark for evaluating the robustness of models trained on larger datasets under realistic long-term and indoor surveillance conditions.
We make the following contributions:
  1. 1.
    We introduce CHIRLA, the first fully annotated, long-term, video-based person Re-ID dataset in an indoor environment, featuring challenging clothing changes, multi-camera views, and higher-resolution bounding boxes than existing datasets. CHIRLA also provides more bounding boxes and more samples per ID than previous clothing-change datasets.
  2. 2.
    We define specific benchmark protocols, not only for person Re-ID but also for person tracking, covering diverse and challenging scenarios such as occlusion, reappearance, and multi-camera transitions.
  3. 3.
    We conduct extensive experiments with CNN-based models [17, 18, 19, 20, 21], Vision Transformers [22, 23, 24, 25, 26], and state-of-the-art tracking methods [27, 28, 29, 30, 31, 32, 33], demonstrating that CHIRLA remains highly challenging and serves as a valuable benchmark for future research.
@@ Methods
This study was conducted under the supervision of the Ethical Committee of the University of Alicante (Approval No. UA-2022-11-12). All participants were informed about the study, including data handling procedures, and provided their consent for participation and unrestricted data use. Specifically, participants granted permission for their inclusion in dataset video recordings, which capture physical data (face and body features), and consented to making these recordings publicly available with unrestricted use. The dataset has been made publicly available under the CC-BY license.
@@
To streamline and expedite the manual verification and correction of bounding boxes and ID annotations, we developed a graphical user interface (GUI). This tool significantly facilitates the labeling process (Figure 3). The GUI is publicly available at https://github.com/bdager/preid-labeling-gui.
@@
Benchmark Data Preparation
After preparing the dataset annotations, we used them to convert the data into a format required for the benchmarks proposed in Section Technical Validation. Different train and test splits were created for each benchmark scenario. Data were initially generated automatically following the procedures described below, after which a manual selection process was applied to determine the final train and test sets.
Tracking Benchmark Data Generation.
For the brief occlusions tracking benchmark, we defined occlusions as ID trajectory interruptions lasting between one and five seconds. In this case, we included for each ID only the frames immediately before and after an occlusion began and ended, along with their corresponding bounding box coordinates. For the multiple-person occlusions scenario, we extended the brief occlusion condition by requiring an intersection between the occluded ID and another ID before the occlusion occurred. We applied an Intersection over Union (IoU) threshold of 0.3 to ensure sufficient bounding box overlap before occlusion. From the generated data, we selected two sequences for training and eight for testing, prioritizing challenging scenarios in both tracking tasks.
Re-Identification Benchmark Data Generation.
For person Re-ID benchmark, we focused on evaluating frames following long occlusions (lasting more than five seconds). This distinction ensures a clear differentiation from tracking scenarios that involve shorter occlusions. In this case, we included in the train sets the first visible frame of each ID, followed by four consecutive frames sampled at an 18-frame gap (equivalent to 2.4 seconds of visibility). By contrast, the test sets comprise the reappearance frames and the four subsequent frames for each reappearance, using the same sampling strategy as the train sets. However, a minimum of 20 frames per ID is required in the test sets. If fewer than 20 frames are available, additional frames are randomly sampled to meet the requirement.
In addition, for IDs present in the dataset, we assigned negative values in the test sets when they meet the long-term occlusion conditions but are not included in the train sets. This labeling explicitly defines them as unknown IDs, what we consider as distractors, providing an additional challenge to the Re-ID system.
In Section Technical Validation, the Re-ID benchmark is divided into four scenarios. For each, we selected subsets of train and test data, with the goal that the final scenario evaluation is computed as the average of individual subset results. All evaluation data was carefully analyzed to align with benchmark definitions, emphasizing challenging cases. Figure 4 illustrates sample data from one Re-ID benchmark scenario.
((a)) Train data samples
((b)) Test data samples
  • Re-identification after reappearance on the scene. We manually selected 10 train and test splits. The train set consists of annotations from 10 videos, each subset containing one video per sequence. However, only nine distinct sequences met the long occlusion condition, requiring the repetition of one sequence with a different sampled video. If uninterrupted frames were unavailable within the analyzed duration (2.4 seconds), only frames within the specified range were included. Each test set included all frames where the individual reappeared after a long occlusion.
  • Long-term re-identification. We selected seven train and test subsets. The train set included annotations from one video per camera, totaling seven videos, with one per subset. The corresponding test set comprised all remaining videos from the same camera analyzed in the train set across different sequences.
  • Multi-camera re-identification. Data was distributed across 10 train and test subsets. The train set consisted of annotations from one video per sequence, totaling 10 videos, with one per subset. The corresponding test subset included all remaining videos from the same sequence.
  • Multi-camera long-term re-identification. Similar to the previous scenario, we selected 10 subsets for training and testing. The train set consisted of annotations from one video per sequence, totaling 10 videos. Each test subset contained two videos from one or two sequences different from those in the train subset.
@@
FNFN ={}={} total number of false negatives over all frames.
FPFP ={}={} total number of false positives over all frames.
IDSWIDSW ={}={} total number of ID switches over all frames.
GTGT ={}={} total number of ground truth detections over all frames.
@@
[…]
@@
By default, precision takes all the retrieved correspondences into account, but it can also be evaluated at a given number of retrieved images, commonly known as cut-off rank, where the model is only assessed by considering its top-most queries. The measure is called precision at $k$ or P@k@k. Using the P@k@k we can calculate the average precision (AP@n@n) for a number $n$ of matches, as shown in Equation 8. The final AP is the value of AP@n@n when all the ground truth positives have been obtained.
@@
GTPGTP ={}={} total number of ground truth positives (correct samples of the query class),
n ={}={} total number of candidate matches ($n$ elements most similar to the query),
$rel@k$ ={}={} relevance function that equals 1 if the ID at rank $k$ is correct.
@@
NN ={}={} total number of queries,
$i$ ={}={} index for the i-th query,
@@
The dataset is composed of a collection of videos, as provided by the set of cameras in the environment, the corresponding Re-ID labels, and train and test data to use in the benchmarks proposed in Section Technical Validation. The Re-ID labels comprise the bounding box of the individuals present in each frame of the video and the corresponding ID. In the following subsections, we will delve into the components of the dataset and how it is organized.
@@
The dataset consists of 10 sequences, each containing seven videos recorded by the seven cameras deployed in the environment. The videos were collected over a period of seven months, allowing the dataset to capture variations in both the environment and the appearance of individuals. The average duration of a single video is 284 seconds, with a range of 113 to 567 seconds. In total, the dataset comprises more than 5.52 hours of video and 596,345 frames, with a total of 963,554 annotated bounding boxes. The details of each sequence are presented in Table 2. The videos are stored in AVI format, encoded using DivX MPEG-4 at 30 fps with a bitrate of 4048 Kbps.
@@
In addition to the video data, we provide Re-ID labels. As mentioned earlier, each video is accompanied by a label file that contains, for each frame, the bounding boxes of all detected individuals along with their corresponding IDs. The labels are stored in JSON format, making them easy to parse. Each JSON file consists of a dictionary where the keys represent the frame numbers of the video, and the corresponding values are lists. Each element in these lists is a dictionary containing two keys:
  • "id": Unique ID assigned to each person.
  • "BboxP": A list of the bounding box coordinates $X_{1},Y_{1},X_{2},Y_{2}]$ that define the region occupied by the person in the analyzed frame.
Frame and ID values are 1-based, whereas bounding box coordinates are 0-based. Specifically, the top-left corner of the bounding box corresponds to $(0,0)$. Figure 5 presents examples of labeled individuals in randomly selected frames from the dataset.
Train and Test Data
Besides the video and Re-ID data, the dataset also includes train and test data specifically selected to use in the benchmarks proposed in Section Technical Validation. For the tracking benchmark, the dataset provides data for two tasks: Tracking after brief occlusions and Tracking in scenarios with multiple people and occlusions. For each scenario, the train set consists of two sequences, while the test set includes eight sequences. In both cases, the data is structured in a hierarchical JSON format. Each JSON file consists of a dictionary where the keys represent individual IDs in the video, and the values are lists of annotations. Each annotation is represented as a dictionary containing four keys:
  • "frame_end": The frame number immediately after the occlusion ends.
For person Re-ID benchmark, same as in tracking benchmark, all data is also organized in a hierarchical JSON structure. Each JSON file consists of a dictionary where the keys represent individual IDs within the video, and the values are lists of annotations. Each annotation is represented as a dictionary with the following two keys:
  • "frame": The frame number where the person is present in the video.
The dataset provides train and test splits for four Re-ID benchmark scenarios. The Re-identification after reappearance on the scene scenario includes 10 train and test splits, each containing one video per sequence, selected from nine distinct sequences. The Long-term re-identification scenario consists of seven subsets, with the train set containing one video per camera and the test set comprising all remaining videos from the same camera. The Multi-camera re-identification scenario is structured into 10 train and test subsets, with each train set including one video per sequence and the corresponding test set containing all remaining videos within the same sequence. Lastly, the Multi-camera long-term re-identification scenario also includes 10 train and test subsets, where each test subset contains two videos from one or two sequences distinct from those in the train set.
@@
The dataset is structured into three primary directories: annotations, videos, and benchmark. The first two directories contain 10 subdirectories corresponding to the individual sequences listed in Table 2. Each sequence is labeled as seq_XXX, where XXX represents a three-digit, zero-padded sequence number (e.g., seq_026). Within each sequence subdirectory, data is further organized by camera views, with seven files per sequence—one per camera. Each file is named based on the camera number, date, and time of recording.
@@
The benchmark directory contains data for benchmarks proposed in Section Technical Validation. This directory is divided into two main subdirectories: tracking and reid, each containing subdirectories for different benchmark scenarios, further split into train and test sets.
The tracking subdirectory consists of two benchmark scenarios: brief_occlusions and multiple_people_ occlusions. Each scenario is divided into train and test subdirectories, organized by sequence name (e.g., seq_004). Within each sequence subdirectory, data is further organized by camera views. Each sequence contains seven JSON files (one per camera), labeled with the camera name and metadata (e.g., camera_7_2023-06-08-12:29:08.json). Alongside these JSON files, an imgs directory is provided, containing seven subdirectories—one per camera view—named according to the corresponding JSON files. Inside each camera view subdirectory, data is structured into subfolders named after the object IDs in the corresponding video. Each ID subfolder contains cropped bounding box images of that ID, selected based on the tracking scenario. These images are stored in PNG format and named as frame_n.png, where n represents the frame number in the original video.
The reid subdirectory contains four re-identification scenarios: reappearance, long_term, multi_camera, and multi_camera_long_term. Each scenario is split into train and test subdirectories, further divided into subsets named train_X or test_X, where X denotes the subset index (e.g., train_0, test_1). The long_term scenario consists of seven subsets in both train and test sets, whereas the other three scenarios (reappearance, multi_camera, and multi_camera_long_term) contain 10 subsets each. The directory structure in the re-identification benchmark follows the same format as in the tracking benchmark, including sequence-based organization, JSON annotation files, and imgs directories.
@@
These tests assess the ability of a system to detect a person in the scene, create an ID for them, and track them on a single camera until the person disappears from the scene for a few seconds. For this purpose, the video sequences are divided into a train set and a test set. The train set is used to optimize certain parameters of the tracking algorithm that make it applicable to this setting, and adapt them to the resolution and the distance at which persons are seen in the scene. On the other hand, the test set will be used to evaluate the performance of the tracking systems, making use of the MOTA, IDF1 and their auxiliary metrics, as well as the generalization capacity of these systems after having optimized their parameters in videos from train sequences only.
With the proposed train and test sets, we check the capabilities of the tracking system in the following scenarios:
@@
For this purpose, the video sequences are divided into train and test sets. The train sets are used to optimize the parameters of the Re-ID algorithm, adapting it to the visual characteristics of the recorded persons, so that it learns to recognize their ID, especially focusing on the characteristics that allow them to be identified in the long term.
On the other hand, the test sets are designed to evaluate the performance of the system in different Re-ID scenarios, using metrics such as mAP, CMC and other auxiliary metrics that allow measuring its generalization capacity after the optimization of parameters in sequences other than train ones.
With the proposed train and test sets, we check the capabilities of the Re-ID system in the following scenarios: